OpenStudy (pphalke):

If $4,300 yields $1,290 in simple interest over six years, what is the annual rate?

OpenStudy (mathstudent55):

If it yields $1290 in 6 years, what does it yield in 1 year?

OpenStudy (mathstudent55):

Simple interest yields the same amount of interest each year. For example, if you deposit $100 at 3% simple interest per year, then each year you earn $3 since $3 is 3% of $100. If you leave the money for 6 years in the bank, you will earn 6 * $3 = $18 in interest over 6 years.

OpenStudy (pphalke):

ok how do I start to solve the problem? I think the answer is 5% I am not sure how to solve

OpenStudy (mathstudent55):

I asked above. If you earn $1290 in 6 years, how much do you earn in 1 year?

OpenStudy (mathstudent55):

Here are two ways of solving this problem: A. 1. First find how much amount of interest is earned in 1 year by dividing the total amount of interest earned by the number of years. 2. Find what percent of $4300 the amount earned in 1 year is. B. 1. Use the formula for simple interest, I = prt, where I = total interest amount earned, p = principal, the amount of money deposited r = annual interest rate t = time in years 2. Plug in the numbers you have into the formula, and solve for i, the interest rate.

OpenStudy (mathstudent55):

Can you follow the methods above?

OpenStudy (dumbcow):

\[1290 = 4300*6*r\] \[r = \frac{1290}{4300*6}\]